# Antonio Raiola

### v1

@@ Abstract
While the literature on grouped patterns in panel data analysis has received significant attention, little to no results are available on testing for their presence. We propose using existing tools for testing slope homogeneity in panels for this purpose. We highlight the key advantages and limitations of the available testing frameworks under a sequence of doubly local alternatives, where slopes are divided into dominant and remainder groups, with the size of the remainder groups and the slopes differences shrinking at a certain rate as the sample size increases. A Monte Carlo study corroborate our theoretical findings.
@@
Keywords: Panel Data; Grouped Patterns; Slope Heterogeneity.
@@ 1 Introduction
Driven by empirical demand, the econometric literature on grouped structure in panel data has received significant attention in both econometrics and economics. Most contributions to date focus on identification, estimation, and inference for individual group membership and group-specific parameters; see, e.g., Bonhomme and Manresa (2015), Bester and Hansen (2016), Su et al. (2016), Lu and Su (2017), and Dzemski and Okui (2024). Before estimating such structures, an applied researcher typically faces a more basic diagnostic question: is there evidence of any grouped structure in the first place? This paper propose a simple and readily implementable solution. We repurpose well-known test for slope homogeneity as diagnostics for the presence of a grouped structure, without estimating group membership or the number of groups. We show that these tests provide exactly the needed diagnostic when evaluated under appropriately designed sequences of local alternatives that encode grouped patterns, and we develop theory that clarifies when and why these tests have nontrivial power.
@@
Our approach is guided by practicality: we rely on well-known tools that are readily available in standard software. The literature on testing slope homogeneity is long-standing. A seminal starting point is Swamy (1970)’s a test statistic, based on deviations of individual slope estimates from the pooled estimator; refinements such as Pesaran and Yamagata (2008) remain among the most widely used tools for panel data models. In parallel, several LM tests have been developed for random-coefficient models, where the null of slope homogeneity is equivalent to the variance of the random coefficients being zero. Building on Breusch and Pagan (1979) (see also Hsiao 2003, Sec. 6.2.2.4), Juhl and Lugovskyy (2014) propose LM statistics for zero variance, and Breitung et al. (2016) further extend the methodology to empirically relevant settings. Other LM tests assess whether the regressors have predictive power for the restricted residuals under the null of slope homogeneity–Su and Chen (2013) in models with interactive fixed effects and Lu and Su (2017), originally developed to determine the number of groups.
@@
We analyze the power of established test statistics—Pesaran and Yamagata (2008), Su and Chen (2013), and the LM-type test of Breitung et al. (2016)—under a sequence of doubly local alternatives designed to capture grouped patterns. These alternatives are doubly local in the sense that (i) the population is split into a dominant group and a remainder set of units (possibly partitioned into several groups), and (ii) the slope differences across groups shrink at a controlled rate. This design allows us to characterize how power depends jointly on the convergence rate of group-specific differences and the relative size of the remainder group(s). We consider this setup across several scenarios: two groups; multiple groups, where the number of groups is either fixed or increases with the cross-sectional dimension $N$; and both large-$N$, large-$T$ panels and large-$N$, fixed-$T$ panels, where $T$ denotes the time dimension.
@@
Our analysis yields several findings. We derive the boundaries that characterize when grouped patterns are detectable by linking the rate at which slope differences vanish, $\gamma$, to the size of the remainder group(s), $M$. In large-$N$, large-$T$ panels, the boundary is governed by $\gamma\asymp M^{1/2}T^{1/2}/N^{1/4}$. This result nests the classical $N^{1/4}\sqrt{T}$ rate—when $M/N\to c>0$—reported in Pesaran and Yamagata (2008), Su and Chen (2013), and Breitung et al. (2016), derived under analogous Pitman-drift sequences. Importantly, the tests remain informative even if $M/N\to 0$, provided $MT\to\infty$ and the boundary condition above is satisfied, making sparse grouped alternatives detectable as well. When $T$ is fixed, detectability requires $M$ to be on the order of $\sqrt{N}$ or larger; below this threshold, grouped heterogeneity is undetectable. Finally, if the number of groups increases with $N$, detectability depends on the total number of units in the remainder groups, rather than on the sheer count of groups. A detailed analysis of each scenario is presented in Section 4. Monte Carlo evidence in Section 5 confirm our theoretical findings.
@@
The rest of the paper is structured as follows. Section 2 introduces the model and notation. Section 3 presents the test statistics. Section 4 develops the local-power analysis under doubly local alternatives for the large-$T$ and fixed-$T$ cases, including multiple-group settings. Proofs are relegated to the appendix.
@@ 2 Preliminaries
We are interested in testing for the presence of a grouped structure in the slopes by means of existing slope homogeneity tests. Specifically, we study the local power of such tests against alternatives in which the slope coefficients differ across groups of individuals. The null hypothesis of identical slopes is formulated as,
@@
As discussed in the introduction, a variety of well-established tests for slope homogeneity are available in the literature. To analyze their asymptotic behavior within a unified framework, we adopt a common set of assumptions sufficiently broad to capture their essential features.
Throughout, $C$ denotes a generic constant that may vary across expressions, and #G\#G denotes the cardinality of a generic set $G\subseteq\mathcal{N}$.
@@
For all $i\in\mathcal{N}$: $(a)$ $\mathbb{E}|x_{it,k}|^{6}<C<\infty$, for all t=1,,Tt=1,...,T and k=1,,Kk=1,...,K; $(b)$ the probability limit $S_{i}=\mathop{\rm plim}_{T\to\infty}T^{-1}x_{i}^{\prime}x_{i}$, is a positive definite, non-stochastic and finite matrix; (c) for any subset $G\subseteq\mathcal{N}$ such that $G\to\infty$, the probability limit $S^{(G)}=\mathop{\rm plim}_{N,T\to\infty}(\#G)^{-1}\sum_{i\in G}S_{i}$ is a positive definite, non-stochastic and finite matrix.
@@ 3 Test Statistics
Under $H_{0}^{**}$ and $N\to\infty$, with $T$ either fixed or $T\to\infty$, the test has a $\chi^{2}_{K}$ null distribution.
The three tests, $\Delta$, $\mathcal{J}$, and $LM$, are consistent for $H_{0}$ and have well-established (local) power properties. However, none has been explicitly studied in the context of local power against grouped slope structures, which is the focus of the next section.
@@ 4 Asymptotic Local Power Analysis
We study the local power of $\Delta$, $\mathcal{J}$, and $LM$ tests under grouped alternatives, focusing on (i) how large slope differences must be to be detected and (ii) how large the group must be for detection. To this end, we introduce a sequence of doubly local alternatives that extends the classical Pitman framework and integrates these two aspects. We first present the results for the case where the sample is partitioned into two groups, $\{G_{1},G_{2}\}$, with $N,T\to\infty$. Group $G_{1}$ is treated as the “dominant” group and $G_{2}$ as the “alternative” one, containing the remaining individuals. We then extend these results to panels with $N\to\infty$ and fixed $T$, as well as to settings with more than two groups.
Formally, we define
$H_{1,n}:\beta_{g(i)}=\begin{cases}\beta,&\text{if }i\in G_{1},\\ \beta+\lambda_{2}/\gamma,&\text{if }i\in G_{2},\end{cases}$
@@
Allowing the scaling factor $\gamma$ to depend on $M$ as well as on $N$ and $T$ reflects that our proposed local alternatives control for both the slope difference and the size of the heterogeneous group. This extension is essential: the relative magnitude of $M$ with respect to $N$ and $T$ is crucial in determining when and why slope homogeneity tests can detect group-based heterogeneity. Theorem 1 formalizes this statement.
@@
${\delta}_{J}=\frac{c^{2}}{\sqrt{V_{0}}}\lambda_{2}^{\prime}\left(W^{(G_{2})}+m_{0}(S^{(G_{2})}S^{-1}WS^{-1}S^{(G_{2})}-2W^{(G_{2})}S^{-1}S^{(G_{2})})\right)\lambda_{2},$
with $W^{(G)}=\mathop{\rm plim}\limits_{\#G,T\to\infty}(\#GT)^{-1}\sum_{i\in G}\sum_{t=1}^{T}{x_{it}x_{it}^{\prime}}(1-h_{i,tt})$, and $W=W^{(\mathcal{N})}$. Here $V_{0}=\mathop{\rm plim}_{N,T\to\infty}4T^{-2}N^{-1}\sum_{i=1}^{N}\sum_{t=2}^{T}\mathbb{E}\left[\left(\epsilon_{it}b_{it}^{\prime}\sum_{s=1}^{t-1}b_{is}\epsilon_{is}\right)^{2}\right]$, with $b_{it}=S_{i}x_{it}$.
@@
Theorem 1 highlights a fundamental trade-off between the rate at which slope differences vanish and the size of the alternative group. Its key implication is the rate
@@
which controls the transition between trivial and nontrivial power. Specifically, if $\gamma$ diverges faster than the rate in (5), all three tests have trivial power; otherwise they have power to detect the grouped alternative.
It is worth emphasizing when the result in (5) reproduces earlier findings and when it yields genuinely new insights. If the fraction of alternative units remains bounded away from zero, $m_{0}>0$ (so that $M\asymp N$), the critical rate in (5) reduces to the classical order $O(\sqrt{T}N^{1/4})$. In this case, our framework nests the local power behavior established by PY, SC, and BRS, providing a consistency check with existing results. The theorem further derives noncentrality parameters that differ from those in the original papers and clarify the distinct role of group-based heterogeneity in local power.
Further novel implications arise when the alternative group is small relative to the dominant one, i.e., when $m_{0}=0$. In this case, the rate in (5) is slower than $O(\sqrt{T}N^{1/4})$; nevertheless, the tests remain informative as long as $M$ diverges with $N$, achieving nontrivial power even when only a small fraction of units (i.e., $M/N\to 0$) display distinct slopes. This contrasts with the conventional intuition in the grouped-patterns literature, where heterogeneous slopes are typically assumed to affect a non-negligible share of the sample (e.g., Bonhomme and Manresa, 2015; Su et al., 2016; Lu and Su, 2017). These results highlight the usefulness of slope homogeneity tests in empirical contexts where the presence of small, systematically different subgroups matters for inference and policy.

### v2

@@ Abstract
Empirical researchers often use slope-homogeneity tests to assess whether slopes can be treated as common across units. A key difficulty is that heterogeneity may be concentrated in a small number of units, so that a failure to reject homogeneity may reflect limited power rather than true homogeneity. We quantify this issue by analyzing the power of standard slope-homogeneity tests under doubly local alternatives — alternatives in which only small groups of units depart from the common slope and the magnitude of the deviations shrinks with sample size. We characterize detectability as a function of panel dimensions, the size of the departing groups, and the rate at which deviations shrink. The results tell the researcher clearly when homogeneity tests are informative and when they will miss small-group heterogeneity. A Monte Carlo study confirms the theory.
@@
Keywords: Panel Data; Detection Boundary; Slope Heterogeneity; Local Power; Group Structures.
@@ 1 Introduction
Slope homogeneity is a convenient and widely used restriction in panel data analysis. If slopes are common across units, pooling information over the cross section improves precision and yields estimators with a simple interpretation. If slopes are heterogeneous, common-slope estimands may be difficult to interpret or even misleading (see, e.g., Breitung and Salish, 2021). Empirical researchers therefore use slope-homogeneity tests to assess whether slopes are common and to choose between pooled and heterogeneous specifications. A key difficulty is that heterogeneity may be concentrated in a small number of units, so that non-rejection may reflect limited power against such departures rather than genuine homogeneity. This paper quantifies the local power of standard slope-homogeneity tests against departures confined to small groups. Our results characterize, in terms of panel dimensions and group size, the conditions under which standard tests have power against such localized departures.
@@
The literature on testing slope homogeneity is long-standing. A seminal starting point is Swamy (1970), based on the dispersion of individual-slope estimates around a pooled estimator; refinements such as Pesaran and Yamagata (2008) remain among the most widely used tools for linear panel models. In parallel, several LM-type procedures have been developed in random-coefficient models, where the null of slope homogeneity is equivalent to a zero-variance restriction on the random coefficients. Building on Breusch and Pagan (1979) (see also Hsiao, 2003, Sec. 6.2.2.4), Juhl and Lugovskyy (2014) propose LM statistics for zero variance, and Breitung et al. (2016) further extend the methodology to empirically relevant settings. Other LM tests assess whether the regressors have predictive power for the restricted residuals under the null of slope homogeneity: Su and Chen (2013) in models with interactive fixed effects, and Lu and Su (2017), whose test for the number of latent groups nests slope homogeneity as a special case.
@@
We analyze the power of established test statistics—Pesaran and Yamagata (2008), Su and Chen (2013) [^2] In the setup considered here, the test of Su and Chen (2013) coincides with that of Lu and Su (2017) when specialized to the single-group null; our results therefore cover both simultaneously., and the LM-type test of Breitung et al. (2016)—under a sequence of doubly local alternatives designed to capture small-group heterogeneity. These alternatives are doubly local in the sense that $(i)$ only the slopes of small groups of individuals deviate from the dominant common slope, and $(ii)$ the slope differences across groups shrink at a controlled (Pitman-type) rate. This design characterizes how power depends jointly on the group-specific slope differences and the size of the groups. We develop the analysis for a benchmark setting with two groups (one dominant majority, one heterogeneous remainder), and for extensions with multiple groups, with either a fixed or growing number of groups. The analysis covers both large-$N$, large-$T$ panels and large-$N$, fixed-$T$ panels, where $T$ denotes the time dimension.
Our main results deliver a detection boundary by linking the rate at which slope differences vanish, $\gamma$, to the size of the heterogeneous group(s), $M$. In large-$N$, large-$T$ panels, this boundary is given by
$\gamma\asymp\frac{M^{1/2}T^{1/2}}{N^{1/4}},$
@@
which nests the classical $N^{1/4}\sqrt{T}$ rate—obtained when $M/N\to c>0$—reported in Pesaran and Yamagata (2008), Su and Chen (2013), and Breitung et al. (2016) under analogous Pitman-drift sequences. Importantly, this boundary shows that the tests remain informative even if $M/N\to 0$, provided $MT\to\infty$ and the boundary condition above holds, thereby enabling detection of small-groups alternatives. When $T$ is fixed, detectability requires $M$ to be of order $\sqrt{N}$ or larger; below this threshold, heterogeneity is undetectable. Finally, when the number of heterogeneous groups increases with $N$, the tests retain nontrivial power and detectability depends on the total number of units outside the dominant majority rather than on the number of groups per se.
The results also speak to the literature on latent group structure in panel models, where such departures arise naturally—see, e.g., Bonhomme and Manresa (2015), Bester and Hansen (2016), Su et al. (2016), Lu and Su (2017), and Dzemski and Okui (2024) for identification, estimation, and inference for group membership and group-specific parameters. Our contribution is complementary, as we characterize when standard slope-homogeneity tests can detect departures consistent with latent group structure. This perspective also connects to recent work on inference when latent groups are not well separated; see Akgun and Okui (2025).
@@
The remainder of the paper is organized as follows. Section 2 introduces the model and notation. Section 3 reviews the test statistics. Section 4 derives the local-power results under doubly local alternatives, and Section 5 reports a Monte Carlo study. Proofs are relegated to the appendix.
@@ 2 Preliminaries
We are interested in testing the presence of group heterogeneity in the slopes by means of existing slope homogeneity tests. The null hypothesis of identical slopes is
@@
As discussed in the introduction, a variety of well-established tests for slope homogeneity are available in the literature. To analyze their asymptotic behavior within a unified framework, we adopt a common set of assumptions sufficiently broad to capture their essential features. Throughout, $C$ denotes a generic constant that may vary across expressions, and #G\#G denotes the cardinality of a generic set $G\subseteq\mathcal{N}$.
@@
For all $i\in\mathcal{N}$: $(a)$ $\mathbb{E}|x_{it,k}|^{6}<C<\infty$, for all t=1,,Tt=1,...,T and k=1,,Kk=1,...,K; $(b)$ the probability limit $S_{i}=\mathop{\rm plim}_{T\to\infty}T^{-1}x_{i}^{\prime}x_{i}$, is a positive definite, non-stochastic and finite matrix; (c) for any subset $G\subseteq\mathcal{N}$ such that $G\to\infty$, the probability limit $S^{(G)}=\mathop{\rm plim}_{N,T\to\infty}(\#GT)^{-1}\sum_{i\in G}x_{i}^{\prime}x_{i}$ is a positive definite, non-stochastic and finite matrix.
@@ 3 Test Statistics
Under $H_{0}^{**}$ and $N\to\infty$, with $T$ either fixed or $T\to\infty$, the test has a $\chi^{2}_{K}$ limiting distribution.
The three tests, $\Delta$, $\mathcal{J}$, and $LM$, are consistent for $H_{0}$ and have well-established local power properties. However, none has been explicitly studied for local power against small-group heterogeneity, which is the focus of the next section.
@@ 4 Asymptotic Local Power Analysis
We study the local power of $\Delta$, $\mathcal{J}$, and $LM$ tests under grouped alternatives, focusing on (i) how large slope differences must be to be detected and (ii) how large the groups must be for detection. To this end, we introduce a sequence of doubly local alternatives that extends the classical Pitman framework and integrates these two aspects. We first present the results for the case where the sample is partitioned into two groups, $\{G_{1},G_{2}\}$, with $N,T\to\infty$. Group $G_{1}$ is treated as the “dominant” group and $G_{2}$ as the “alternative” one, containing the remaining individuals. We then extend these results to panels with $N\to\infty$ and fixed $T$, as well as to settings with more than two groups.
Formally, define
$H_{1,n}:\beta_{g(i)}=\begin{cases}\beta,&\text{if }g(i)=G_{1},\\ \beta+\lambda_{2}/\gamma,&\text{if }g(i)=G_{2},\end{cases}$
@@
Allowing the scaling factor $\gamma$ to depend on $M$ as well as on $N$ and $T$ reflects that our proposed local alternatives control for both the slope difference and the size of the heterogeneous group. This extension is essential: the relative magnitude of $M$ with respect to $N$ and $T$ determines when slope homogeneity tests can detect group-based heterogeneity. Theorem 1 formalizes this statement.
@@
${\delta}_{\mathcal{J}}=\frac{c^{2}}{\sqrt{V_{0}}}\lambda_{2}^{\prime}\left(W^{(G_{2})}+m_{0}(S^{(G_{2})}S^{-1}WS^{-1}S^{(G_{2})}-2W^{(G_{2})}S^{-1}S^{(G_{2})})\right)\lambda_{2},$
with $W^{(G)}=\mathop{\rm plim}\limits_{\#G,T\to\infty}(\#GT)^{-1}\sum_{i\in G}\sum_{t=1}^{T}{x_{it}x_{it}^{\prime}}(1-h_{i,tt})$, and $W=W^{(\mathcal{N})}$. Here $V_{0}=\lim_{N,T\to\infty}4T^{-2}N^{-1}\sum_{i=1}^{N}\sum_{t=2}^{T}\mathbb{E}\left[\left(\epsilon_{it}b_{it}^{\prime}\sum_{s=1}^{t-1}b_{is}\epsilon_{is}\right)^{2}\right]$, with $b_{it}=S^{-1/2}_{i}x_{it}$.
@@
Theorem 1 highlights a trade-off between the rate at which slope differences vanish and the size of the alternative group. The critical rate,
@@
marks the boundary between trivial and nontrivial power. If $\gamma$ grows faster than this rate, all three tests have trivial power. If it grows at or below this rate, they have power to detect the grouped alternative.
Two cases deserve attention. If the fraction of alternative units remains bounded away from zero ($m_{0}>0$, so $M\asymp N$), the rate (5) reduces to the classical order $O(\sqrt{T}N^{1/4})$, consistent with the local power results of PY, SC, and BRS. When the alternative group is small ($m_{0}=0$), the rate is slower than $O(\sqrt{T}N^{1/4})$, yet the tests retain nontrivial power as long as $M$ diverges with $N$ — even as the fraction of units with distinct slopes vanishes. This contrasts with the prevailing assumption in the grouped-patterns literature, which requires heterogeneous slopes to affect a non-negligible share of the sample (e.g., Bonhomme and Manresa, 2015; Su et al., 2016; Lu and Su, 2017).
On the noncentrality parameters.
Theorem 1 shows that all three tests share the same asymptotic detection boundary (5). Their local power need not coincide, however, since the noncentrality parameters differ. A general comparison is intractable, as the parameters depend on both model and design in complex ways. We therefore consider the benchmark case of a single regressor, $K=1$, to build intuition.
Let $x_{it}$ be scalar and i.i.d. across $i$ and $t$, with $\mathbb{E}\left[x_{it}\right]=0$, $\mathbb{E}\left[x_{it}^{2}\right]=\sigma_{i,x}^{2}$, and uniformly bounded fourth moments. The regressor may thus exhibit unit-specific variance, constant over time. For any subset $G\subseteq\mathcal{N}$, define
$\mu_{2,x}^{(G)}=\lim_{\#G\to\infty}\frac{1}{\#G}\sum_{i\in G}\sigma_{i,x}^{2},\quad\mu_{4,x}^{(G)}=\lim_{\#G\to\infty}\frac{1}{\#G}\sum_{i\in G}\bigl(\sigma_{i,x}^{2}\bigr)^{2},\quad q_{x}^{(G)}=\lim_{\#G\to\infty}\frac{1}{\#G}\sum_{i\in G}\frac{\sigma_{i,x}^{2}}{\sigma_{i}^{2}},$
with $\mu_{2,x}=\mu_{2,x}^{(\mathcal{N})}$, $\mu_{4,x}=\mu_{4,x}^{(\mathcal{N})}$, and $q_{x}=q_{x}^{(\mathcal{N})}$. Now suppose $m_{0}=0$, so $M/N\to 0$ and the heterogeneous subset is asymptotically negligible. The second term in each noncentrality parameter vanishes, and Theorem 1 reduces to
$\delta_{\Delta}=\frac{c^{2}\lambda_{2}^{2}}{\sqrt{2}}\,q_{x}^{(G_{2})},\qquad\delta_{\mathcal{J}}=\frac{c^{2}\lambda_{2}^{2}}{\sqrt{V_{0}}}\,\mu_{2,x}^{(G_{2})},$
and
$\xi_{LM}=\delta_{LM}^{\prime}\Psi^{-1}\delta_{LM}=\frac{c^{4}\lambda_{2}^{4}}{4\Psi}\,\left(\mu_{4,x}^{(G_{2})}\right)^{2},$
where $V_{0}=2\lim_{N\to\infty}N^{-1}\sum_{i=1}^{N}\sigma_{i}^{4}$ and $\Psi=\lim_{N\to\infty}(2N)^{-1}\sum_{i=1}^{N}\sigma_{i}^{4}(\sigma_{i,x}^{2})^{2}$. Even in the scalar case, PY and SC depend on different weighted second moments, while BRS depends on fourth-order cross-sectional moments of the regressor.
To compare PY and SC directly, note that
$\frac{\delta_{\Delta}}{\delta_{\mathcal{J}}}=\lim_{N\to\infty}\mathrm{RMS}(\sigma^{2}_{i})\sum_{i\in G_{2}}\frac{\omega_{x,i}}{\sigma^{2}_{i}}\;$
where $\omega_{x,i}=\sigma^{2}_{x,i}/\sum_{i\in G_{2}}\sigma^{2}_{x,i}$, and $\mathrm{RMS}(\sigma^{2}_{i})=\sqrt{N^{-1}\sum_{i=1}^{n}\sigma^{4}_{i}}$. It follows that,
$\lim_{N\to\infty}\frac{\mathrm{RMS}(\sigma^{2}_{i})}{\max_{i\in G_{2}}\sigma^{2}_{i}}\leq\frac{\delta_{\Delta}}{\delta_{\mathcal{J}}}\leq\lim_{N\to\infty}\frac{\mathrm{RMS}(\sigma^{2}_{i})}{\min_{i\in G_{2}}\sigma^{2}_{i}}.$
When $G_{2}$ concentrates on units with small $\sigma^{2}_{i}$, the $\Delta$ test dominates; otherwise the $\mathcal{J}$ test has higher local power. Under homoskedasticity ($\sigma_{i}^{2}=\sigma^{2}$ for all $i$), $\delta_{\Delta}=\delta_{\mathcal{J}}$ and the two tests are locally equivalent.
Comparing PY and SC with BRS is more complex. Under homoskedasticity and constant regressor variance ($\sigma_{i,x}^{2}=\sigma_{x}^{2}$ for all $i$), all three procedures are locally equivalent: $\xi_{LM}/\delta_{\Delta}^{2}=\xi_{LM}/\delta_{\mathcal{J}}^{2}=1$. If errors are homoskedastic and $\{\sigma_{x,i}^{2}\}$ unrestricted with $\mu_{2,x}^{(G_{2})}=\mu_{2,x}$, and $\mu_{4,x}^{(G_{2})}=\mu_{4,x}$, then
$\frac{\xi_{LM}}{\delta_{\Delta}^{2}}=\frac{\mu_{4,x}}{\mu_{2,x}^{2}}\geq 1.$
Hence, BRS has a larger local noncentrality parameter than PY and SC, with the ratio $\xi_{LM}/\delta_{\Delta}^{2}$ governed by the cross-sectional dispersion in regressor variance.
A complementary comparison is obtained when regressor variance is constant across units but errors are heteroskedastic. Then, $\delta_{\mathcal{J}}^{2}=\xi_{LM}$, so SC and BRS tests are locally equivalent. If $G_{2}$ is also asymptotically representative with respect to error variances — that is $\lim_{M\to\infty}M^{-1}\sum_{i\in G_{2}}\sigma_{i}^{-2}=\lim_{N\to\infty}N^{-1}\sum_{i=1}^{n}\sigma_{i}^{-2}$ — then
$\frac{\xi_{LM}}{\delta_{\Delta}^{2}}=\left(\frac{\delta_{\mathcal{J}}}{\delta_{\Delta}}\right)^{2}=\lim_{N\to\infty}\frac{1}{\left(N^{-1}\sum_{i=1}^{N}\sigma_{i}^{4}\right)\left(N^{-1}\sum_{i=1}^{N}\sigma_{i}^{-2}\right)^{2}}\leq 1,$
with equality under homoskedasticity. When regressor variance is constant, error heteroskedasticity thus favors PY over SC and BRS
When $m_{0}>0$ (i.e., $M\asymp N$), the comparison is more complex as the second term in each expression of noncentrality parameters is no longer asymptotically negligible. However, the analysis of the same scalar benchmark points in the same direction. In particular, under homoskedasticity, $\delta_{\Delta}=\delta_{\mathcal{J}}$; if, in addition, the regressor variance is constant, all three tests are locally equivalent; if, instead, regressor variance is unrestricted and $\mu_{2,x}^{(G_{2})}=\mu_{2,x}$, and $\mu_{4,x}^{(G_{2})}=\mu_{4,x}$,
$\frac{\xi_{LM}}{\delta_{\Delta}^{2}}=\frac{\mu_{4,x}}{\mu_{2,x}^{2}}\geq 1.$
Outside these benchmarks, the comparison depends on the joint distribution of $\{\sigma^{2}_{i}\}$ and $\{\sigma^{2}_{x,i}\}$ across the full sample and the remainder group $G_{2}$.